OpenStudy (australopithecus):

Let f(x) = e^(x^(3)). Calculate the 9th derivative f^(9)(0) using the exponential series and Taylor’s formula.

OpenStudy (kainui):

What part are you needing help with? You surely know how to find every derivative up to the 9th, right? If not, we can talk about that or talk about what taylor's formula is if that's where you're having difficulty. Taylor and Mclauren series just mean that you're making a polynomial function that has the same derivatives at a point, up to as many derivatives as you take.

OpenStudy (australopithecus):

yes I know how to find there derivative up to the 9th. I just don't know how to answer this

OpenStudy (anonymous):

9!/3!

OpenStudy (anonymous):

\[ \large e^{x^3}= \sum_{n=0}^\infty \frac{x^{3n}}{n!}\] Now I would expand it